OpenStudy (anonymous):

Evaluate the expression, log base 8^163840 - log base 8^5?

OpenStudy (anonymous):

Try to Factorize 163840 in terms of 8 and 5

OpenStudy (anonymous):

163840 = 8^5 * 5

OpenStudy (anonymous):

\[\log_{8} (8^5*5) = \log_{8}8^5 + \log_{8}5 \]

OpenStudy (anonymous):

\[\log_{8}8^5 + \log_{8}5 - \log_{8}5 = \log_{8}8^5 \]

OpenStudy (anonymous):

\[=\frac{ 5 }{1 } = 5\]

OpenStudy (anonymous):

Ahh I finally understand it, thank you very much!!

OpenStudy (anonymous):

Welcome...)